Dublin University Rifle Club

Version Date : May 16, 2001 Filename : Status :

Dublin University Rifle Club

Running a successful Trinity Top Ten

Mark Dennehy

May 16, 2001

I collected the notes in this document from running the Trinity Top Ten ( T.T.T) for two years. Since I had to learn the job on the fly at the start, I felt that a set of notes was the best legacy I could leave to whoever had to be the Match Secretary after me; I hope they prove useful to you.

Mark Dennehy

DURC Match Secretary 1999-2002

• Be organised early on. I can’t stress this enough - if you fall behind in marking cards, you will not catch up and will have to duck people at competition as they ask about the results and who’s won and so forth. So get yourself organised well ahead of time.

• Set the dates before you send out the first invitation. Ensure everyone knows whether the date is when the card is to be shot by; or posted by; or with you by.

• Ensure people know that dropping 2 cards will result in disqualification.

• Be prepared to spend one saturday each fortnight or each month (depending on the length of time for shooting each card) taking all the cards for that round and scoring them. Scoring them as they arrive would be better, but that’s like doing your homework before watching TV...

• Bring home the .22 gauge for these sessions. You do know how to use it, right ? Also have a well-lit area with a desk and a magnifying glass for this.

• When scoring the cards, keep a written record of the scores in a notebook and archive the cards themselves. Record as well who has submitted cards for what round and so on. This makes it a lot easier if you have to reenter all the results into a computer because you accidentally overwrote the full set of results with an earlier incomplete backup ...

• Spreadsheets a mystery? Then learn them. Now. You really cannot do this job without the aid of a computer – well, actually you can, but you can’t do anything else at the same time because it will eat all your time!

• Unix, L^{A}TEX, unix spreadsheets like gnumeric or kspread, and plain ASCII text file formats (comma seperated value – .csv files) are all your friends. Get to know them and your life will be a lot easier than you can currently imagine. The ability to do work over a modem line from home is one which you will find yourself in dire need of at least once during your year. That’s Murphy’s Law for you.

• Buy a hardbacked notebook and keep notes on everything. Computers are wonderful tools that mean you can do this job without giving up your BA/MSc/PhD; but the hard disk you can bet your life on hasn’t been invented yet.

• The first invites usually go out at the DURC 50m Open. Post them off the day before and bring copies to the range for the open to hand out. Give a month to six weeks for replies, then send off the stickers.

• Once the entry fees are in, buy the prizes immediately. In fact, hanging onto the entry fees until the prizes are bought and paid for is a better idea than releasing them to the mercies of the front arch locker – the first year I did that and the entire entry fee income was spent in one night on beer after the fencing club grudge match.

• Remember, ten patches for the top ten shooters, and a medal for the top three.

Acronyms

T.T.T Trinity Top Ten